# Orion: A Fully Homomorphic Encryption Framework for Deep Learning

### v2

@@ Abstract.
Fully Homomorphic Encryption (FHE) has the potential to substantially improve privacy and security by enabling computation directly on encrypted data. This is especially true with deep learning, as today, many popular user services are powered by neural networks in the cloud. One of the major challenges facing wide-scale deployment of FHE-secured neural inference is effectively mapping these networks to FHE primitives. FHE poses many programming challenges including packing large vectors, automatically managing noise via bootstrapping, and translating arbitrary and general-purpose programs to the limited instruction set provided by FHE. These challenges make building large FHE neural networks intractable using the tools available today.
In this paper we address these challenges with Orion, a fully-automated framework for private neural inference in FHE. Orion accepts deep neural networks written in PyTorch and translates them into efficient FHE programs. We achieve this by proposing a novel single-shot multiplexed packing strategy for arbitrary convolutions and through a new, efficient technique to automate bootstrap placement. We evaluate Orion on common benchmarks used by the FHE deep learning community and outperform state-of-the-art by 2.38×2.38$\times$2.38 × on ResNet-$20$, the largest network they report. Orion extends naturally to larger networks. We demonstrate this by evaluating ResNet-50 on ImageNet and present the first high-resolution homomorphic object detection experiments using a YOLO-v1 model with $139$ million parameters. Finally, we open-source our framework Orion at the following repository: https://github.com/baahl-nyu/orion.
**footnotetext: Equal contribution. 🖂 Austin Ebel is the corresponding author.
@@ 1. Introduction
Fully Homomorphic Encryption (FHE) is a powerful encryption scheme that allows for computation to be performed directly on encrypted data without ever needing to decrypt. FHE provides a solution to the privacy concerns of outsourced cloud computation by allowing personal and sensitive data to be first encrypted under FHE before being processed securely in the cloud. In this manner, both the original data and any intermediate results remain encrypted and can only be observed in the clear when the data owner decrypts it locally using their secret key. FHE has broad implications for areas such as finance or health and other situations involving sensitive data. Many of these services are now driven by deep learning, specifically neural networks.
@@
Despite this, the wide-scale deployment of FHE-enabled private deep learning remains limited because of 1) the significant computational costs of FHE and 2) the programming challenges of translating standard, unencrypted neural networks into FHE programs. While hardware accelerators have addressed this performance gap (reagen, ; feldmann2021f1, ; Kim_2022, ; CraterLake, ; soni2023rpu, ; sharp, ; osiris, ) and FHE compilers have made progress in translating workloads into FHE programs  (porcupine, ; coyote, ; CHET, ; EVA, ; viand2023heco, ; fhelipe, ), it remains a challenge to run FHE neural inference on large datasets in a high-level deep learning framework (e.g., ImageNet inference in PyTorch). This paper focuses on efficiently and automatically mapping neural networks, as implemented in modern deep learning libraries, to FHE programs. Multiple compounding factors make effectively mapping neural networks to FHE challenging.
@@
Data packing: FHE encrypts large vectors that vary in length from $2^{14}$ to $2^{17}$, with recent work favoring the larger lengths (soni2023rpu, ; Kim_2022, ; dacapo, ). Scalar data must be packed into the slots of these vectors and FHE only allows SIMD addition, SIMD multiplication, and cyclic rotation upon these encrypted vectors. Individual encrypted elements cannot be indexed, and any function applied to a vector is applied to all its elements. Therefore, it is crucial to optimize data packing and layout in FHE vectors to efficiently make use of the available slots and leverage this SIMD property.
@@
Bootstrap placement: Each encrypted vector can only execute a finite number of computations before decryption fails. To prevent this, bootstrapping can be used to reset this computational budget but comes at an extremely high latency cost (DHBSGS, ). Effectively inserting bootstrap operations into an FHE program is challenging; bootstrapping at any point in an FHE circuit directly affects the latency and resource requirements of subsequent operations, including future bootstraps. Therefore, an efficient solution to bootstrap placement requires a holistic understanding of the entire FHE program.
@@
Programmability: FHE is notoriously hard to program. Every program must be translated into a series of SIMD additions, multiplications, and cyclic rotations on very large vectors. This translation is especially challenging for deep learning because 1) convolutional and fully-connected layers must be implemented using just these three FHE operations and 2) element-wise nonlinear activation functions must be approximated with high-degree polynomials. Furthermore, most existing FHE libraries are not designed with machine learning in mind and are cumbersome for building large FHE programs. Finally, several auxiliary FHE operations (e.g. rescaling, bootstrapping, level adjustment) must be manually inserted into FHE programs making it hard for researchers and practitioners to immediately make use of FHE.
@@
To address these challenges, we develop and open-source Orion: a fully-automated framework for private neural inference that advances the state-of-the-art in all three areas. For data packing, we introduce a novel packing strategy for convolutions that we call single-shot multiplexing. This approach directly improves upon the multiplexed packing technique of Lee et al. (lee2022, ) in three ways: (1) it halves the multiplicative depth of every convolution from two to one, (2) it supports convolutions with arbitrary parameters (e.g., stride, padding, groups, dilation), and (3) it reduces (expensive) ciphertext rotations in modern neural networks by up to 6.41×6.41$\times$6.41 ×.
@@
We then propose a fully-automated and efficient solution for placing bootstrap operations within a neural network. Our approach uses a linear time topological sort to determine the optimal placement of bootstraps. For example, it takes Orion just $6.67$ ($49.4$) seconds to place bootstraps in ResNet-20 (ResNet-110) and requires no user input. Our algorithm scales linearly with network depth and results in a similar number of bootstraps to Dacapo (dacapo, ) while also being up to 255×255$\times$255 × faster. Notably, our method also places just $351$ bootstrap operations in a ResNet-50 evaluated on ImageNet, whereas $8480$ bootstraps are placed in HeLayers (helayers2, ). For ResNet-20, Fhelipe (fhelipe, ) performs 58 bootstraps, whereas we have only 37.
@@
Finally, we design Orion to be accessible to machine learning researchers and practitioners who are unfamiliar with FHE. Orion inherits directly from PyTorch and extends the functionality of its most popular CNN layers. For instance, users can train Orion networks with existing PyTorch training scripts and directly load the weights of pre-trained models with torchvision. Orion’s interoperability with PyTorch allows us to easily verify that our FHE outputs match the outputs of running inference directly in PyTorch.
@@
We evaluate Orion using a set of neural networks and datasets and report significant improvements over state-of-the-art (lee2022, ; dacapo, ; fhelipe, ). To the best of our knowledge, we are the first to make the following contributions:
@@
    We develop and open-source Orion: a fully-automated framework for running FHE neural network inference directly in PyTorch. Using Orion on the standard ResNet-20 FHE benchmark, we achieve a speedup over state-of-the-art (Fhelipe (fhelipe, )) by 2.38×2.38$\times$2.38 ×. We also run ResNet-50 on ImageNet with no code changes, which most prior work does not readily support.
@@
    We implement the single-shot multiplexed packing strategy in Orion and support arbitrary convolutions while only having a multiplicative depth of one.
@@
    We implement our automatic bootstrap placement algorithm in Orion, which requires no user input, and matches state-of-the-art in terms of number of bootstraps while performing up to 255×255$\times$255 × faster.
@@
    We are the first to perform high-resolution object detection using a YOLO-v1 model with $139$ million parameters on an image of size $448\times 448\times 3$ (yolov1, ).
@@ 2. CKKS Background
In this section, we describe the CKKS homomorphic encryption scheme that is used in Orion. At a high level, CKKS encrypts large vectors of complex (or real) numbers and enables three operations on these vectors: element-wise addition, element-wise multiplication, and cyclic rotation (ckks, ). These properties make CKKS a natural choice for applications that operate on real-valued vectors such as deep learning. Below, we detail the fundamental datatypes and operations in CKKS.
@@ 2.1. Datatypes
The three datatypes in CKKS are cleartexts, plaintexts, and ciphertexts. A cleartext is an unencrypted vector of complex (or real) numbers. The process of encoding converts a cleartext into a plaintext, which is still unencrypted and is an element of the polynomial ring $\mathcal{R}_{Q}=\mathbb{Z}_{Q}[X]/(X^{N}+1)$. Here, $N$ is a power of two and each coefficient is an integer modulo $Q$. In this way, a plaintext is a polynomial up to degree $N-1$ and each coefficient is in $\mathbb{Z}_{Q}$.
A plaintext can then be encrypted into a ciphertext which consists of two polynomials (i.e., an element in $\mathcal{R}_{Q}\times\mathcal{R}_{Q}$). To maintain 128-bit security and enable deep arithmetic computations in CKKS, it is common for $N$ to be between $2^{14}$ to $2^{17}$ and $Q$ to be on the order of hundreds to thousands of bits  (cryptoeprint:2019/939, ). In practice, these constraints mean plaintexts and ciphertexts are KBs to MBs in size. CKKS performs polynomial operations on both plaintexts and ciphertexts that correspond to SIMD addition, SIMD multiplication, and cyclic rotations on the underlying cleartext vectors.
@@ 2.2. Encoding: Cleartext $\rightarrow$ Plaintext
Encoding converts a cleartext vector $\mathbf{m}\in\mathbb{C}^{N/2}$ into a plaintext polynomial $[\mathbf{m}]\in\mathcal{R}_{Q}$ by 1) applying an inverse Fast Fourier Transform (iFFT) on the cleartext values, 2) multiplying each output by a scaling factor $\Delta$, and 3) rounding each element to the nearest integer. In this manner, one can pack $n=N/2$ complex values into a single plaintext. If the underlying data consists only of real numbers (such as our use case), it is possible to pack $n=N$ real values into a single plaintext. Decoding converts a plaintext polynomial back into a cleartext vector by performing the FFT on the polynomial coefficients and dividing each value by the scaling factor $\Delta$. Other encoding and decoding schemes for CKKS exist, but do not preserve the SIMD property on the underlying cleartext values (see Section IV.A of  (Kim2023, )).
@@ 2.3. Encryption: Plaintext $\rightarrow$ Ciphertext
A plaintext $[\mathbf{m}]\in\mathcal{R}_{Q}$ is encrypted into a ciphertext $[\![\mathbf{m}]\!]\in\mathcal{R}_{Q}\times\mathcal{R}_{Q}$ by adding noise $[\mathbf{e}]\in\mathcal{R}_{Q}$ to the plaintext and encrypting via either the secret or public key. Decryption can only occur via the secret key. Both rounding during encoding and the addition of noise during encryption introduce small errors, and so CKKS is an approximate homomorphic encryption scheme. However, these approximations tend to be tolerable for deep learning.
@@ 2.4. RNS-CKKS
Each coefficient of a CKKS polynomial can be up to thousands of bits wide (e.g. $\log_{2}Q\approx 1500$), making operations on these large coefficients compute intensive. Therefore, polynomials are decomposed into $L+1$ residual polynomials, each with small coefficients that fit within a machine-sized word. Now, the coefficients of a CKKS polynomial can be viewed as matrix of size $(L+1)\times N$ where each row of the matrix is an $N$-degree residual polynomial with small coefficients.
This process occurs via the Residual Number System (RNS) in which the large modulus $Q$ is decomposed into $L+1$ smaller moduli $Q=\prod_{i=0}^{L}q_{i}$. Using RNS, a coefficient $x\in\mathbb{Z}_{Q}$ can be represented using $L+1$ limbs as $x:=(x_{0},\ldots,x_{L})$ where $x_{i}=x\bmod q_{i}$. Modular addition (multiplication) between two large elements in $\mathbb{Z}_{Q}$ is equivalent to performing modular addition (multiplication) on each limb independently  (Garner1959, ).
The integer $L$ is known as the maximum multiplicative level of the polynomial, and as we will see, homomorphic multiplications consume levels. At any given time, a ciphertext can be at level $\ell$, such that $0\leq\ell\leq L$, and a ciphertext at level $\ell$ consists of $\ell+1$ residual polynomials. When a ciphertext reaches level $\ell=0$, it has depleted its multiplicative budget. A bootstrap procedure can be used to increase its level up to $L_{\text{eff}}$ to perform further multiplications.

### v3

@@ Abstract.
Fully Homomorphic Encryption (FHE) has the potential to substantially improve privacy and security by enabling computation directly on encrypted data. This is especially true with deep learning, as today, many popular user services are powered by neural networks in the cloud. Beyond its well-known high computational costs, one of the major challenges facing wide-scale deployment of FHE-secured neural inference is effectively mapping these networks to FHE primitives. FHE poses many programming challenges including packing large vectors, managing accumulated noise, and translating arbitrary and general-purpose programs to the limited instruction set provided by FHE. These challenges make building large FHE neural networks intractable using the tools available today.
In this paper we address these challenges with Orion, a fully-automated framework for private neural inference using FHE. Orion accepts deep neural networks written in PyTorch and translates them into efficient FHE programs. We achieve this by proposing a novel single-shot multiplexed packing strategy for arbitrary convolutions and through a new, efficient technique to automate bootstrap placement and scale management. We evaluate Orion on common benchmarks used by the FHE deep learning community and outperform state-of-the-art by $2.38\times$ on ResNet-$20$, the largest network they report. Orion’s techniques enable processing much deeper and larger networks. We demonstrate this by evaluating ResNet-50 on ImageNet and present the first high-resolution FHE object detection experiments using a YOLO-v1 model with $139$ million parameters. Orion is open-source for all to use at: https://github.com/baahl-nyu/orion.
fully homomorphic encryption, compilers, cryptography, privacy-preserving machine learning
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 30-April 3, 2025; Rotterdam, Netherlandsbooktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’25), March 30-April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3676641.3716008isbn: 979-8-4007-1079-7/2025/03ccs: Security and privacy Cryptographyccs: Software and its engineering Compilers**footnotetext: Equal contribution. 🖂 Austin Ebel is the corresponding author.
@@ 1. Introduction
Fully Homomorphic Encryption (FHE) is a powerful encryption scheme that allows for computation to be performed directly on encrypted data without ever needing to decrypt. In this manner, both the original data and any intermediate results remain encrypted and can only be observed in the clear when the data owner decrypts it locally using their secret key. FHE provides a solution to the security and privacy concerns of outsourced cloud computation. This has broad implications for areas such as finance or health and other situations involving sensitive data. Many of these services are now driven by deep learning, specifically neural networks.
@@
Today, the wide-scale deployment of FHE-enabled private deep learning remains limited due to the significant computational costs of FHE and the programming challenges of translating standard, unencrypted neural networks into FHE programs. While hardware accelerators have addressed this performance gap (Reagen et al., 2021; Feldmann et al., 2021; Kim et al., 2022; Samardzic et al., 2022; Soni et al., 2023; Kim et al., 2023; Ebel and Reagen, 2024) and FHE compilers have made progress in translating workloads into FHE programs  (Cowan et al., 2021; Malik et al., 2023; Dathathri et al., 2019, 2020; Viand et al., 2023; Krastev et al., 2024), it remains a challenge to run FHE neural inference on large datasets in a high-level deep learning framework (e.g., ImageNet inference in PyTorch). This paper focuses on efficiently and automatically mapping neural networks, as implemented in modern deep learning libraries, to FHE programs. Besides the computational costs, multiple compounding factors make effectively mapping neural networks to FHE challenging.
@@
Vector packing: The CKKS FHE scheme (Cheon et al., 2016, 2018) encrypts large vectors that typically vary in length from $2^{14}$ to $2^{17}$, with recent work favoring the longer lengths (Soni et al., 2023; Kim et al., 2022; Cheon et al., 2024). Scalar data must be packed into vector slots, as CKKS supports only SIMD addition, SIMD multiplication, and cyclic rotation on these encrypted vectors. As a result, performance is directly related to how densely one can fill these slots with meaningful data. Individual encrypted slots cannot be efficiently indexed and any function applied to a vector is applied to all its elements. Therefore, it is crucial to optimize data packing and layout in CKKS vectors to efficiently make use of the available slots and leverage this SIMD property.
@@
Bootstrap placement: Each encrypted vector can only execute a finite number of computations before decryption fails. To address this shortcoming, bootstrapping can be used to reset the computational budget but comes at an extremely high latency cost (Bossuat et al., 2020). Effectively inserting bootstrap operations into an FHE program is challenging; bootstrapping at any point in an FHE circuit directly affects the latency of subsequent operations as well as the placement of future bootstraps. Therefore, an efficient solution to bootstrap placement requires a holistic understanding of the FHE program.
@@
Programmability: Existing FHE libraries are not designed with machine learning in mind and are cumbersome for building large FHE programs. Every program must be translated into a series of SIMD additions, SIMD multiplications, and cyclic rotations on very large vectors. This translation is especially challenging for deep learning since encrypted convolutional and fully-connected layers cannot be evaluated using traditional tensor arithmetic, and because element-wise nonlinear activation functions must be approximated with high-degree polynomials. Finally, several auxiliary FHE operations (e.g., rescaling, bootstrapping, and level adjustment) must be manually inserted into FHE programs making it hard for practitioners to immediately make use of FHE.
@@
To address these challenges, we develop and open-source Orion: a fully-automated framework for private neural inference that advances the state-of-the-art in all three areas. For data packing, we introduce a novel packing strategy for convolutions that we call single-shot multiplexing. This approach directly improves upon the multiplexed packing technique of Lee et al. (Lee et al., 2022b) in three ways: (i) it halves the multiplicative depth of every strided convolution from two to one, (ii) it supports convolutions with arbitrary parameters (e.g., stride, padding, groups, and dilation), and (iii) it reduces expensive ciphertext rotations in modern networks by up to $6.41\times$.
@@
We then present a fully-automated and efficient solution for placing bootstrap operations within a neural network, which we express as a directed acylic graph (DAG). Our approach finds the shortest paths (i.e., latencies) through related DAGs to determine the locations of bootstrap operations that abide by the constraints of FHE. Our algorithm scales linearly with network depth and matches state-of-the-art (Cheon et al., 2024) in terms of number of bootstraps, while also being between $8\times$ and $1270\times$ faster. As an example, it takes Orion just $1.94$ ($11.0$) seconds to place bootstraps in ResNet-20 (ResNet-110) and requires no user input. Notably, our method also places just $351$ bootstrap operations in a ResNet-50 evaluated on ImageNet, whereas $8480$ bootstraps are placed in HeLayers (Baruch et al., 2023). For ResNet-20, Fhelipe (Krastev et al., 2024) (PLDI ’24) performs $58$ bootstraps, whereas we have only $37$.
@@
Finally, we design Orion to be accessible to machine learning researchers and practitioners who are unfamiliar with FHE. Orion inherits directly from PyTorch and extends the functionality of its most popular CNN layers. For instance, users can train Orion networks with existing PyTorch training scripts and directly load the weights of pre-trained models with torchvision. Orion’s interoperability with PyTorch allows us to easily verify that our FHE outputs match the outputs of running inference directly in PyTorch. We demonstrate our capabilities by running, to the best of our knowledge, the largest FHE inference to date: YOLO-v1 (Redmon et al., 2016) on the PASCAL-VOC dataset (Everingham et al., 2010).
@@
We evaluate Orion using a set of neural networks and datasets and report significant improvements over state-of-the-art (Lee et al., 2022b; Cheon et al., 2024; Krastev et al., 2024). We make the following contributions:
@@
    We develop and open-source Orion: a fully-automated framework for running FHE neural network inference directly in PyTorch. Using Orion on the standard ResNet-20 FHE benchmark, we achieve a speedup over state-of-the-art (Fhelipe (Krastev et al., 2024)) by $2.38\times$.
@@
    We present the single-shot multiplexed packing strategy in Orion and support arbitrary convolutions while only having a multiplicative depth of one.
@@
    We implement our automatic bootstrap placement algorithm in Orion, which requires no user input, and matches the state-of-the-art (Cheon et al., 2024) in terms of number of bootstraps while running up to $1270\times$ faster.
@@
    We are the first to perform high-resolution object detection using a YOLO-v1 model with $139$ million parameters on an image of size $448\times 448\times 3$ (Redmon et al., 2016); to this best of our knowledge, this is the largest FHE computation to date.
@@ 2. CKKS Background
In this section, we describe the CKKS homomorphic encryption scheme used in Orion. At a high level, CKKS encrypts large vectors of complex (or real) numbers and supports three operations: element-wise addition, element-wise multiplication, and cyclic rotation (Cheon et al., 2016). These properties make CKKS a natural choice for applications that operate on real-valued vectors such as deep learning. Below, we detail the fundamental datatypes and operations in CKKS. Please refer to Table 1 for our notation.
@@ 2.1. Datatypes
The three datatypes in CKKS are cleartexts, plaintexts, and ciphertexts. A cleartext is an unencrypted vector of complex (or real) numbers. Encoding converts a cleartext into a plaintext, which is still unencrypted and is an element of the polynomial ring $\mathcal{R}_{Q}=\mathbb{Z}_{Q}[X]/(X^{N}+1)$. Here, $N$ is a power of two and each coefficient is an integer modulo $Q$. For technical reasons, the modulus $Q$ is typically chosen to be a product of $L+1$ primes $q_{i}$ such that $q_{i}\bmod 2N=1$ and $\log_{2}q_{i}\approx w$ where $w$ is a standard machine-word size (32 or 64) (Cheon et al., 2018). In this way, a plaintext is a polynomial up to degree N1N-1 and each coefficient is in $\mathbb{Z}_{Q}$.
Table 1. Relevant CKKS parameters.
Param. Description
$\mathbf{m}$ A cleartext vector of real or complex numbers.
$\mathbf{m}]$ A plaintext polynomial encoding $\mathbf{m}$.
$\mathbf{m}]\!]$ A ciphertext encrypting the plaintext $\mathbf{m}]$.
NN Power-of-two polynomial ring degree.
$n$ Length of the vector message, $n\leq N$ (or $\nicefrac{{N}}{{2}}$).
$L$ Maximum multiplicative level of a ciphertext.
$\ell$ Current multiplicative level.
$Q$ Initial polynomial modulus.
$q_{i}$ Small moduli in RNS decomposition of $Q=\prod_{i=0}^{L}q_{i}$.
$\Delta$ Scaling factor for $\mathbf{m}]$ and $\mathbf{m}]\!]$.
$L_{\text{boot}}$ Number of levels consumed by bootstrapping.
$L_{\text{eff}}$ Maximum achievable level after bootstrapping.
A plaintext can then be encrypted into a ciphertext that consists of two polynomials (i.e., an element in $\mathcal{R}_{Q}\times\mathcal{R}_{Q}$). To maintain 128-bit security and enable deep arithmetic computations in CKKS, it is common for $N$ to be between $2^{14}$ to $2^{17}$ and $Q$ to be on the order of hundreds to thousands of bits  (Albrecht et al., 2019). In practice, these constraints mean plaintexts and ciphertexts are KBs to MBs in size. CKKS performs polynomial operations on both plaintexts and ciphertexts that correspond to SIMD complex addition, SIMD complex multiplication, and cyclic rotations on the underlying cleartext vectors. For our use case, we encrypt real-valued vectors rather than complex-valued vectors and therefore utilize encrypted SIMD real addition and SIMD real multiplication.
@@ 2.2. Encoding: Cleartext $\rightarrow$ Plaintext
Encoding converts a cleartext vector $\mathbf{m}\in\mathbb{C}^{N/2}$ into a plaintext polynomial $\mathbf{m}]\in\mathcal{R}_{Q}$ by (i) applying an inverse Fast Fourier Transform (iFFT) on the cleartext values, (ii) multiplying each output by a scaling factor $\Delta$, and (iii) rounding each element to the nearest integer. In this manner, one can pack $n=N/2$ complex values into a single plaintext. If the underlying data consists only of real numbers (such as our use case), it is possible to pack $n=N$ real values into a single plaintext. Decoding converts a plaintext polynomial back into a cleartext vector by performing the FFT on the polynomial coefficients and dividing each value by the scaling factor $\Delta$. Other encoding and decoding schemes for CKKS exist, but do not preserve the SIMD property on the underlying cleartext values (see Section IV.A of  (Kim and Guyot, 2023)).
@@ 2.3. Encryption: Plaintext $\rightarrow$ Ciphertext
A plaintext $\mathbf{m}]\in\mathcal{R}_{Q}$ is encrypted into a ciphertext $\mathbf{m}]\!]\in\mathcal{R}_{Q}\times\mathcal{R}_{Q}$ by adding noise $\mathbf{e}]\in\mathcal{R}_{Q}$ to the plaintext and encrypting via either the secret or public key. Decryption can only occur via the secret key. Both rounding during encoding and the addition of noise during encryption introduce small errors, and so CKKS is an approximate homomorphic encryption scheme. However, these approximations tend to be tolerable for deep learning.
(a) $\mathsf{PMult}$
(b) $\mathsf{HRot}$
(c) $\mathsf{Bootstrap}$
Figure 1. The latencies of key homomorphic operations as a function of ciphertext level ($\ell$) for ring degree $N=2^{16}$ and $\Delta\approx 2^{40}$. Setting $L_{\text{eff}}$ too low would require many low-latency bootstraps, while setting it too high would result in fewer but higher-latency bootstraps. We set $L_{\text{eff}}=10$ in order to balance bootstrap latency with having a reasonable number of remaining levels for computation.
@@ 2.4. RNS-CKKS
The coefficients of a CKKS polynomial can be up to thousands of bits wide (e.g., $\log_{2}Q\approx 1500$), making operations on these large coefficients compute intensive. Therefore, polynomials are typically decomposed into $L+1$ residual polynomials, each with small coefficients that fit within a machine-sized word. In this way, each coefficient in a CKKS polynomial is represented as a vector of length $L+1$, and it follows that the $N$ coefficients for a single CKKS polynomial can be organized as a matrix of size $(L+1)\times N$. Here, each row of the matrix is an $N$-degree residual polynomial with small coefficients (if all $q_{i}$ are chosen appropriately, then every element of the matrix will fit within a single 32-bit or 64-bit machine word).
This process occurs via the Residual Number System (RNS) in which the large modulus $Q$ is decomposed into $L+1$ smaller moduli $Q=\prod_{i=0}^{L}q_{i}$. Using RNS, a coefficient $x\in\mathbb{Z}_{Q}$ can be represented using $L+1$ limbs as $x:=(x_{0},\ldots,x_{L})$ where $x_{i}=x\bmod q_{i}$. Modular addition (multiplication) between two large elements in $\mathbb{Z}_{Q}$ is equivalent to performing modular addition (multiplication) on each limb independently  (Garner, 1959).
The integer $L$ is known as the maximum multiplicative level of the polynomial, and as we will see, homomorphic multiplications consume levels. At any given time, a ciphertext at level $\ell$ actually has coefficients taken not from $\mathbb{Z}_{Q}$ but rather from $\mathbb{Z}_{Q_{\ell}}$ where $Q_{\ell}=\prod_{i=0}^{\ell}q_{i}$. As a result, the coefficients of a CKKS polynomial at level $\ell$ are actually represented as a matrix of size $(\ell+1)\times N$ rather than $(L+1)\times N$. For this reason, homomorphic operations at lower levels are less costly. The latency of key homomorphic operations at various levels can be seen in Figure 1. When a ciphertext reaches level $\ell=0$, it has depleted its multiplicative budget, and a bootstrap procedure can be used to increase its level up to $L_{\text{eff}}$ to perform further multiplications. For the interested reader, we note that the superlinear increase in bootstrapping latency observed in Figure 1c results from an increase in the decomposition number (dnum) when key-switching to maintain $128$-bit security (Han and Ki, 2020; Neda et al., 2024).